# Authorship Verification based on the Likelihood Ratio of Grammar Models

### v1

@@ 1.5 Language Models and Compression Methods
In addition to IM and its variants, the most successful family of methods in AV are the ones that are based on language modeling and/or compression, which are mathematically related. A language model is here defined as a generative statistical model that assigns a probability distribution to stretches of language.
One of the very first successful examples for this class of methods was the winner of the PAN 2015 competition, Bagnall [8], whose AV method was based on a character-level recurrent neural network language model. Another notable neural network language modeling approach is Boenninghoff’s et al. [14] use of a Siamese neural network. Their approach was found to outperform IM by halving its error rate while also offering a way to visualize the importance of tokens. Their approach was designed to mitigate the effect of content but its limitation is that a large and structured comparison dataset is necessary, which is not always available. King and Cook [61] also proposed an approach based on language models which works by retraining the final layer of a model based on potentially short samples from an author and then comparing this model against a general model on the questioned text. They found that this system outperformed the use of traditional machine learning methods like Support Vector Machines. Although language models are also inevitably affected by the issue of content bias, attempts to mitigate this problem are scarce. A notable mention is Fourkioti’s et al. [37] approach consisting in training an $n$-gram language model on sequences of POS tags. The authors found that this kind of model is only valuable when fused with other features.
@@
Related to language modeling, the other approach that has been found to be very successful is the use of compression algorithms. This family of methods is fundamentally very similar to language models because they both are probabilistic models of a training dataset. Language modeling can be thought of as a process of compression and one of the best compressor for text data, prediction by partial matching (PPM), could also be seen as a character-level language model. A particularly successful AV method which makes use of PPM as an underlying engine is COAV [46, 45]. The method was found to perform similarly to IM, but in a much shorter time and without relying on impostors. Halvani’s 2021 survey of the field [41] indeed concluded that COAV, together with IM, are the best AV methods. Both of them can be used together with POSNoise to avoid content bias. Although IM tended to be more robust over the various tests it was subjected to, the advantage of COAV over IM is that the method is deterministic and much faster (for comparison see, for example, [46, Table 4]). In addition, Halvani [41] was able to visualize the importance of character sequences for the prediction by producing a color-coded version of the texts, thus making the method more qualitatively explorable by the analyst.
@@ 2 The Likelihood Ratio of Grammar Models
The method proposed in this paper is a binary-extrinsic AV method influenced by the literature above which offers a general theoretically-motivated framework for AV that can help to better explain the other methods reviewed above. This section explains its theoretical framework and the algorithm, which is outlined in Algorithm 1, while most technical details are left to Section 5.
Input: $\mathbb{D}_{\mathcal{U}}=\{\mathcal{D}_{\mathcal{U}}^{1},\mathcal{D}_{\mathcal{U}}^{2},\dots\}$, $\mathbb{D}_{\mathcal{A}}=\{\mathcal{D}_{\mathcal{A}}^{1},\mathcal{D}_{\mathcal{A}}^{2},\dots\}$, $\mathbb{D}_{\text{ref}}=\{\mathcal{D}_{\text{ref}}^{1},\mathcal{D}_{\text{ref}}^{2},\dots\}$, $N$, $r$
1 Function sample($k$, $\mathbb{S}$):
2       Randomly sample a subset of $k$ sentences from the input set $\mathbb{S}$
3Function sent($\mathbb{D}$):
4       Construct a set $\mathbb{S}$ of tokenized sentences from the input set $\mathbb{D}$
// Apply POSNoise to each document set
$\mathbb{S}_{\mathcal{U}}\mbox{}\leftarrow\textnormal{{sent}}(\textsf{POSNoise}\mbox{}(\mathbb{D}_{\mathcal{U}}))$ $\mathbb{S}_{\mathcal{A}}\mbox{}\leftarrow\textnormal{{sent}}(\textsf{POSNoise}\mbox{}(\mathbb{D}_{\mathcal{A}}))$ $\mathbb{S}_{\textsf{ref}}\mbox{}\leftarrow\textnormal{{sent}}(\textsf{POSNoise}\mbox{}(\mathbb{D}_{\text{ref}}))$ // Build Grammar Model for the known author $\mathcal{A}$
$G_{\mathcal{A}}\leftarrow\textnormal{{KN}}(\mathbb{S}_{\mathcal{A}},N)$ // Build reference Grammar Models
5 for $i\leftarrow 1$ to $r$ do
6       $\mathbb{S}_{\text{i}}\leftarrow$ sample($|\mathbb{S}_{\mathcal{A}}|$, $\mathbb{S}_{\text{ref}}$) $G_{i}\leftarrow\textnormal{{KN}}(\mathbb{S}_{\text{i}},N)$
@@ 4       Construct a set $\mathbb{S}$ of tokenized sentences from the input set $\mathbb{D}$
7 $\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})\leftarrow 0$ for $S\mbox{}\in\mathbb{S}_{\mathcal{U}}$ do
       // Decompose $S\mbox{}$ into a sequence of tokens
8       $(t\mbox{}_{1},t\mbox{}_{2},\ldots,t\mbox{}_{m})\leftarrow S\mbox{}$ for $j\leftarrow 1$ to $m$ do
             // Calculate mean $\lambda_{G}$ for $t\mbox{}_{j}$ over reference Grammar Models
9             for $i\leftarrow 1$ to $r$ do
10                   $c\leftarrow(t\mbox{}_{j-N+1},\ldots,t\mbox{}_{j-1})$ $\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})\leftarrow\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})+\frac{1}{r}\log\frac{P(t\mbox{}_{j}|c;G_{\mathcal{A}})}{P(t\mbox{}_{j}|c;G_{i})}$
11
12
return $\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})$
@@ 8       $(t\mbox{}_{1},t\mbox{}_{2},\ldots,t\mbox{}_{m})\leftarrow S\mbox{}$ for $j\leftarrow 1$ to $m$ do
[…]
$P(S\mbox{};G)=\prod_{j=1}^{m}P(t\mbox{}_{j}|t\mbox{}_{j-N+1},t\mbox{}_{j-N+2},\ldots,t\mbox{}_{j-1};G)$ (1)
@@
For readability reasons, we rewrite the probability assigned by $G$ to each token $t\mbox{}$ in its context $c$ as $P(t\mbox{}|c;G)$ (as in line 18 in Algorithm 1, where $t\mbox{}_{j}$ is used instead of $t$). We estimate Grammar Models adopting Kneser-Ney smoothing [62, 26], represented in Algorithm 1 as $\textnormal{{KN}}()$, with a fixed single default discount parameter $D=0.75$, as explained in Section 5.3. We set $N=10$ as an order of the model to ensure that all grammatical relations among the tokens of a sentence are captured, even in cases of long sentences. However, the order of the model $N$ could be considered as a hyperparameter of the algorithm (lines 8 and 11 in Algorithm 1).
@@
Assuming the Principle of Linguistic Individuality that at any time $\tau$, for any language $L_{\tau}$, there do not exist two individuals, $\mathcal{A}$ and $\mathcal{B}$, for whom their grammar is identical [82], we expect that in a large number of cases the two Grammar Models $G_{\mathcal{A}}$ and $G_{\mathcal{B}}$ would assign two different probabilities to the same token in context. $P(t\mbox{}|c;G_{\mathcal{A}})$ can be interpreted as the extent to which token $t\mbox{}$ is ’grammatical’ or, according to Cognitive Linguistic theories, entrenched for author $\mathcal{A}$ when in context $c$. For this reason, we would expect that, if the sequence $s=t\mbox{}_{1},t\mbox{}_{2},t\mbox{}_{3}$ was produced by $G_{A}$, then $P(s;G_{\mathcal{A}})>P(s;G_{\mathcal{B}})$ or, more generally according to the Principle of Linguistic Individuality, $P(s;G_{\mathcal{A}})>P(s;G_{\neg\mathcal{A}})$. The model $G_{\mathcal{A}}$ is estimated from $\mathbb{S}_{\mathcal{A}}\mbox{}$ (line 8 in Algorithm 1). Instead, to generate an approximation to $G_{\neg\mathcal{A}}$ we estimate a set of reference Grammar Models $\mathbb{S}_{\textsf{ref}}\mbox{}=\{G_{1},G_{2},\dots,G_{r}\}$ where each $G_{r}$ is generated by uniformly sampling randomly from $\mathbb{S}_{\textsf{ref}}\mbox{}$ a set of sentences $\mathbb{S}_{r}$ of the same size as the set of sentences in the candidate author’s corpus, $\mathbb{S}_{\mathcal{A}}\mbox{}$, so that $|\mathbb{S}_{r}|=|\mathbb{S}_{\mathcal{A}}\mbox{}|$ (line 10 in Algorithm 1). We set $r=100$ but this value can also be considered a hyperparameter of the algorithm. A comparison of hyperparameter settings is reported in Section 3.
Taking an approach consistent with the likelihood ratio framework adopted in forensic science, similarly to Ishihara [48], it is therefore possible to calculate (lines 16-18 in Algorithm 1) a mean log-likelihood ratio of the probability assigned to each token in context by $G_{\mathcal{A}}$ and by each $G_{r}$, which we call $\lambda_{G}$:
$\lambda_{G}(t\mbox{}|c)=\frac{1}{r}\sum_{i=1}^{r}{\log\frac{P(t\mbox{}|c;G_{\mathcal{A}})}{P(t\mbox{}|c;G_{i})}}$ (2)
@@
In Eq. 2 the numerator can be seen as a measure of similarity, how much the token is grammatical or entrenched for $G_{\mathcal{A}}$ given its context, while the denominator is a measure of typicality, how ’ungrammatical’ or atypical this token in context is in the reference population. A value of $\lambda_{G}$ for each $S\mbox{}\in\mathbb{S}_{\mathcal{U}}$ can then be calculated (line 14-18 in Algorithm 1) as:
$\lambda_{G}(S\mbox{})=\sum_{j=1}^{m}\lambda_{G}(t\mbox{}_{j}|c)$ (3)
@@
Since we model each $S\mbox{}\in\mathbb{S}_{\mathcal{U}}\mbox{}$ as independent from each other, we can then calculate (line 12-18 in Algorithm 1):
$\begin{split}\lambda_{G}(\mathcal{D}_{\mathcal{U}})=\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})=\sum_{S\mbox{}\in\,\mathbb{S}_{\mathcal{U}}\mbox{}}\lambda_{G}(S\mbox{})=\sum\limits_{S\mbox{}\,\in\,\mathbb{S}_{\mathcal{U}}\mbox{}}\sum\limits_{j=1}^{m}{\frac{1}{r}\sum\limits_{i=1}^{r}\log\frac{P(t\mbox{}_{j}|c;G_{\mathcal{A}})}{P(t\mbox{}_{j}|c;G_{i})}}\end{split}$ (4)
@@
An advantage of this formula is that the final $\lambda_{G}$ can be decomposed into $\lambda_{G}$ values for each single sentence and each single token in each single sentence. This means that the analyst can reconstruct step by step what features contribute to the final score, for example, by ranking each sentence for importance and, for each sentence, color-code or highlight the most important tokens.

### v2

@@ 1.4.3 Language Models and Compression Methods
In addition to IM and its variants, the most successful family of methods in AV are the ones that are based on language modeling and/or compression, which are mathematically related. Here, a language model is defined as a generative statistical model that assigns a probability distribution to stretches of language.
One of the very first successful examples for this class of methods was the winner of the PAN 2015 competition, Bagnall [12], whose AV method was based on a character-level recurrent neural network language model. Another noteworthy neural language modeling approach, based on a Siamese neural network, is HRSN  [^4] HRSN stands for Hierarchical Recurrent Siamese Network, proposed by Boenninghoff et al. [20]. This approach was found to outperform IM by reducing its error rate by half. However, as noted by the authors [19], it has two key drawbacks: first, it does not utilize linguistically interpretable stylometric features, and second, it lacks a mechanism for directly visualizing the decision-making process. Additionally, HRSN is constrained by its reliance on a large and well-structured comparison dataset, which may not always be available.
Another well-known AV method that can be regarded as an improved version of HRSN is ADHOMINEM  [^5] ADHOMINEM stands for Attention-based Deep Hierarchical cOnvolutional siaMese bIdirectional recurreNt nEural-network Model [19]., which was also proposed by Boenninghoff et al. [19]. ADHOMINEM achieved the highest ranking in the PAN AV competition 2020 [70] and, according to Lundmark et al. [87], has shown great promise with respect to real AV applications. The method can be summarized as follows: Given a verification case $c=(\mathcal{D}_{\mathcal{U}},\mathcal{D}_{\mathcal{A}})$, ADHOMINEM preprocesses both documents by hierarchically breaking them down into sentences, words, and characters. A Siamese neural network, incorporating a CNN and bi-directional LSTM layers with attention mechanisms, is then employed to derive a neural document embedding for each input. Specifically, the CNN layer extracts word representations from character embeddings, which are subsequently passed to an LSTM layer operating at the sentence level. The final LSTM layer synthesizes these sentence embeddings to construct a holistic document representation. Based on the document embeddings of $\mathcal{D}_{\mathcal{U}}$ and $\mathcal{D}_{\mathcal{A}}$, a nonlinear metric learning scheme is employed, which measures how closely $\mathcal{D}_{\mathcal{U}}$ and $\mathcal{D}_{\mathcal{A}}$ resemble each other.
Although language models are also inevitably affected by the issue of content bias, attempts to mitigate this problem are scarce. A notable mention is Fourkioti’s et al. [46] approach consisting in training an $n$-gram language model on sequences of POS tags. The authors found that this kind of model is only valuable when fused with other features.
@@
Related to language modeling, the other approach that has been found to be very successful is the use of compression algorithms. This family of methods is fundamentally very similar to language models because they both are probabilistic models of a training dataset. Language modeling can be thought of as a process of compression and one of the best compressor for text data, Prediction by Partial Matching (PPM), could also be seen as a character-level language model. A particularly successful AV method which makes use of PPM as an underlying engine is COAV [57, 56]. The method was found to perform similarly to IM, but in a much shorter time and without relying on impostors. Halvani’s 2021 survey of the field [51] indeed concluded that COAV, together with IM, are the best AV methods. Both of them can be used together with POSNoise to avoid content bias. Although IM tended to be more robust over the various tests it was subjected to, the advantage of COAV over IM is that the method is deterministic and much faster (for comparison see, for example, [57, Table 4]). Given a verification case $c=(\mathcal{D}_{\mathcal{U}},\mathcal{D}_{\mathcal{A}})$, COAV compresses $\mathcal{D}_{\mathcal{U}}$, $\mathcal{D}_{\mathcal{A}}$ and the concatenation of both using a PPM variant called PPMd. Based on the length of the three compression results, a dissimilarity value is calculate using Compression-based Cosine (CBC). This dissimilarity is then compared to a threshold that is optimized on a training corpus to estimate whether either $\mathcal{A}=\mathcal{U}$ or $\mathcal{A}\neq\mathcal{U}$ holds. In addition, Halvani [51] was able to visualize the importance of character sequences for the prediction by producing a color-coded version of the texts, thus making the method more qualitatively explorable by the analyst.
1.4.4 Feature Vector Difference
Finally, another AV method that is also worth mentioning, and which was ranked second in the PAN AV competition 2020 [70], is the so-called Feature Vector Difference approach proposed by Weerasinghe and Greenstadt [132]. Given a verification case $c=(\mathcal{D}_{\mathcal{U}},\mathcal{D}_{\mathcal{A}})$, the first step of this method is to construct corresponding feature vectors $\mathcal{F}\mbox{}_{\mathcal{U}},\mathcal{F}\mbox{}_{\mathcal{A}}$ using a set of features. Then, based on these two, a difference vector $\mathcal{F}\mbox{}=|\mathcal{F}\mbox{}_{\mathcal{U}}-\mathcal{F}\mbox{}_{\mathcal{A}}|$ is then calculated, which represents either a Y- or N-case. This procedure is performed for all verification cases in a training corpus, and based on these a classifier is finally trained. For this, Weerasinghe and Greenstadt considered two classification models: Logistic Regression (LR) and Multi-Layer Perceptron (MLP). Regarding the features, the authors chose tf-idf values of: character $n$-grams (for $n\in\{1,2,\ldots,6\}$), POS tag trigrams, 31 predefined special characters, trigrams of POS tag chunks (extracted from the parse tree representation of the document) as well as noun and verb phrase expansions. Moreover, they considered frequencies of 179 stopwords (defined in the NLTK corpus package [^6] https://github.com/nltk/nltk), total number of characters and tokens in the document, average word length (characters per word) per document, distribution of token lengths (relative frequency of tokens with length $\ell$ (for $1\leq\ell\leq 10$) and vocabulary richness measures.
2 Cognitive Linguistics, Grammar, and Linguistic Individuality
After reviewing the AV literature, this section introduces the theoretical grounding for the new method that we are proposing, which is motivated and inspired by the field of Cognitive Linguistics.
In Cognitive Linguistics, language and grammar are seen as complex adaptive systems [13], where regularities emerge from speakers interacting with each other, like a phenomenon of the third kind [40]. This is partially evidenced by decades of linguistics studies on corpora which have revealed that the traditional understanding of grammar as a set of phrase structure rules acting over basic Part-of-Speech categories (e. g., nouns, adjectives) is untenable and that, instead, a better model of grammar is a lexicogrammar continuum that goes from fixed lexicalized expressions (e. g., “I don’t know”) to flexible templates (e. g., “the” X-er “the” Y-er) up to completely abstract structures (e. g., Agent Action Patient).
This continuum roughly corresponds to the continuum between declarative memory, which is semantic/encyclopedic memory for facts and which tends to be conscious, and procedural memory, which is instead memory for the probabilistic prediction of patterns of sequences and that deals with subconscious knowledge of habits and skills [128]. The roles that these two memories play in language is clearly evidenced in cases of aphasia. When an impairment affects circuits of declarative memory, this gives rise to fluent aphasia, resulting in speech that is mostly constituted by function words (articles, prepositions, pronouns) or morphological forms (“-ed”, “-s”) but with lack of content words that contain meaning, leading to fluent but nonsensical speech, almost resembling POSNoise pre-processed data such as the one in Table 7. Instead, when the impairment is on procedural memory, then speech is only consisting of unconnected content words [129]. These two memory systems work together and, typically, once a habit is formed through repetition then procedural memory takes over, leading to fluency of expression. This transition explains why languages tend to be heavily formulaic. For example, in English one can say “big mistake” but not “large mistake”, even though a formal phrase structure rule based grammar would seemingly have no problems with the latter expression, which is an adjective modifying a noun. What explains this reliance on formulaic expressions is the cognitive tendency to rely as much as possible on habitual and ritualized patterns stored as templates [128, 115].
The way these units of procedural memory are formed is through the process of chunking: when two items that often go together are found together, then these two items can be treated as one, thus saving space in working memory [48, 27, 34]. In all effects, chunking is a process of information compression and, if the units of grammar are chunks, then grammar itself can be thought of as a compressed version of a language or, equivalently, as a probability model over these units. According to Cognitive Linguistics, therefore, a probability model like a language model is a better model of grammar than a list of phrase structure rules.
This way of conceiving language also explains why each individual has a different and possibly unique grammar, a statement encapsulated by the Principle of Linguistic Individuality [97]. The status of unithood given by chunking is not discrete (e. g., something is or is not a unit) but a continuum of strength of unithood, where this strength is called entrenchment [39, 111]. Because entrenchment is dependent on regular exposure and usage, it is fundamentally personal and idiosyncratic. It is for this reason that Langacker hypothesized that “the set of units mastered by all members of a speech community might be a rather small proportion of the units constituting the linguistic ability of any given speaker” [85], a hypothesis that is finding increasing empirical support [97]. For example, for most if not all speakers of English the entrenchment for a sequence such as “of the” is maximal, meaning that they can produce it effortlessly and unconsciously as a single unit. In contrast, a sequence such as “the of” would not be a unit for speakers of English because it is virtually never encountered or used. If a speaker of English encounters this sequence, they would decompose it as a structure made up of two units, “the” and “of”. In between these two extreme examples there is a lot of freedom as to what counts as a unit for a person or another, depending on what each individual has automatized in their procedural memory. For example, an individual $\mathcal{A}_{1}$ might have a very entrenched sequence such as “it is known that” when writing academic articles while an individual $\mathcal{A}_{2}$ might have “it is well established that”. The higher the entrenchment, the more likely it is that $\mathcal{A}$ is using procedural memory and that it is therefore performing an unconscious routine that constitutes a unit for this individual. These sequences will then have open slots ready to be filled in with the appropriate meaning.
Although there have been other attempts at modeling the grammar of an author in the past for AV or AA purposes [31, 11, 46, 50], these methods relied too much on a conception of grammar in the traditional sense, for example using sequences of POS tags or syntactic trees. For the reasons outlined above, these ways of modeling grammar are not compatible with predictions made in Cognitive Science and Cognitive Linguistics. The picture of grammar painted by Cognitive Linguistics is much more lexically-based and word-idiosyncratic. For example, the preposition “of” in English behaves very oddly compared to other prepositions [115]. Giving it the same POS tag as to other prepositions is therefore going to obscure a lot of important grammatical information.
With this theoretical grounding, it is now possible to present our proposed novel approach, LambdaG.
@@ 3 The Likelihood Ratio of Grammar Models
The algorithm presented in this paper is a binary-extrinsic AV method. This section provides an overview of the algorithm, which is formally outlined in Algorithm 1 and schematically illustrated in Figure 1. For a more detailed technical discussion, refer to Section 6.
Figure 1: Schematic overview of our proposed AV method LambdaG.
Input: $\mathcal{D}_{\mathcal{U}}$, $\mathcal{D}_{\mathcal{A}}$, $\mathbb{D}_{\text{ref}}=\{\mathcal{D}_{\text{ref}}^{1},\mathcal{D}_{\text{ref}}^{2},\dots\}$, $N$, $r$
// where $\mathcal{D}_{\mathcal{U}}$, $\mathcal{D}_{\mathcal{A}}$ are two documents, $\mathbb{D}_{\text{ref}}=\{\mathcal{D}_{\text{ref}}^{1},\mathcal{D}_{\text{ref}}^{2},\dots\}$ is a set of reference documents, $N$ is the model order and $r$ is the number of repetitions
1
2Function sample($s$, $\mathbb{S}$):
3       Randomly select a subset of $s$ sentences from the given tokenized input sentence set $\mathbb{S}$
4
5
6Function posnoise($\mathbb{D}$):
7       Applies POSNoise to each document in $\mathbb{D}$ and returns a set of their posnoise representations
8
9Function sent($\mathbb{D}$):
10       Construct a set $\mathbb{S}$ of tokenized sentences from all documents in $\mathbb{D}$
11
// Apply POSNoise to all respective documents and construct tokenized sentences
12 $\mathbb{S}_{\mathcal{U}}\mbox{}\leftarrow\textnormal{{sent}}(\textnormal{{posnoise}}(\{\mathcal{D}_{\mathcal{U}}\}))$
13 $\mathbb{S}_{\mathcal{A}}\mbox{}\leftarrow\textnormal{{sent}}(\textnormal{{posnoise}}(\{\mathcal{D}_{\mathcal{A}}\}))$
14 $\mathbb{S}_{\textsf{ref}}\mbox{}\leftarrow\textnormal{{sent}}(\textnormal{{posnoise}}(\mathbb{D}_{\text{ref}}))$
15
// Build Grammar Model for the known author $\mathcal{A}$
16 $G_{\mathcal{A}}\leftarrow\textnormal{{KN}}(\mathbb{S}_{\mathcal{A}},N)$
17
// Build reference Grammar Models
18 for $j\leftarrow 1$ to $r$ do
19       $\mathbb{S}_{\text{j}}\leftarrow$ sample($|\mathbb{S}_{\mathcal{A}}|$, $\mathbb{S}_{\text{ref}}$)
20       $G_{j}\leftarrow\textnormal{{KN}}(\mathbb{S}_{\text{j}},N)$
21
22
@@ 20       $G_{j}\leftarrow\textnormal{{KN}}(\mathbb{S}_{\text{j}},N)$
23 $\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})\leftarrow 0$
24
25for $S\mbox{}_{i}\in\mathbb{S}_{\mathcal{U}}$ do
       // Decompose $S\mbox{}_{i}$ into a sequence of tokens
26       $(t\mbox{}_{1},t\mbox{}_{2},\ldots,t\mbox{}_{z})\leftarrow S\mbox{}_{i}$
27       for $k\leftarrow 1$ to $z$ do
             // Calculate mean $\lambda_{G}$ for $t\mbox{}_{k}$ over reference Grammar Models
28             for $j\leftarrow 1$ to $r$ do
29                   $\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})\leftarrow\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})+\dfrac{1}{r}\log\dfrac{P(t\mbox{}_{k}|t_{<k};G_{\mathcal{A}})}{P(t\mbox{}_{k}|t_{<k};G_{j})}$
30
31
32
33 return $\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})$
@@ 26       $(t\mbox{}_{1},t\mbox{}_{2},\ldots,t\mbox{}_{z})\leftarrow S\mbox{}_{i}$
We define the most basic grammatical unit of analysis as a function token $t\mbox{}\in\mathcal{T}\mbox{}_{L}$, where $\mathcal{T}\mbox{}_{L}$ is the set all function tokens of a language $L$, including all function words, punctuation marks as well as abstract grammatical categories. For example, for English, $\mathcal{T}\mbox{}_{E}=\{\texttt{the},\texttt{of},\texttt{at},\dots,\allowbreak\texttt{NOUN},\allowbreak\texttt{VERB},\allowbreak\texttt{ADJECTIVE},\dots,\texttt{!},\texttt{:},\texttt{;}\dots\}$. We then define a Grammar Model $G$ as a statistical model that generates a probability distribution over sentences $\mathbb{S}$, which are defined as complete sequences of function tokens. To obtain these representations, we first apply the POSNoise algorithm (see Section 6.2) to $\mathcal{D}_{\mathcal{U}}$, $\mathcal{D}_{\mathcal{A}}$, and each $\mathcal{D}$ in the reference corpus $\mathbb{D}_{\text{ref}}$, retaining only function tokens. This is indicated in Algorithm 1, lines 7–9, where POSNoise is applied. Then, we tokenize each $\mathcal{D}$ into sentences by adding sentence boundaries markers at an end of sentence punctuation mark or new line, thus turning them into their respective set of sentences $\mathbb{S}_{\mathcal{A}}\mbox{}$, $\mathbb{S}_{\mathcal{U}}\mbox{}$ and $\mathbb{S}_{\textsf{ref}}\mbox{}$ (also lines 7–9 in Algorithm 1). The boundaries of each sentence could also be set by a grammatical parsing algorithm. However, we observed that LambdaG is robust to this variation. We then employ $n$-gram models, which allow the probability of any grammatical token to depend only on a finite number of preceding tokens from the same sentence, thereby capturing only the information contained in short-term grammatical correlations. $n$-gram models assume that each sentence is independent and we maintain that this is not an unreasonable assumption since grammatical dependencies between tokens are known to be relatively local [34]. The probability assigned to each sentence $S\mbox{}_{i}=(t\mbox{}_{1},t\mbox{}_{2},\ldots,t\mbox{}_{z})$ by a Grammar Model $G$ would then be calculated as follows:
$P(S\mbox{}_{i};G)=\prod_{k=1}^{z}P(t\mbox{}_{k}|\underbrace{t\mbox{}_{k-N+1},t\mbox{}_{k-N+2},\ldots,t\mbox{}_{k-1}}_{t\mbox{}_{<k}};G)$ (1)
@@
For readability reasons, we rewrite the probability assigned by $G$ to each token $t\mbox{}_{k}$ in its context as $P(t\mbox{}_{k}|t\mbox{}_{<k};G)$ (as in line 19 in Algorithm 1). We estimate Grammar Models adopting Kneser-Ney smoothing [76, 32], represented in Algorithm 1 as $\textnormal{{KN}}()$, with a fixed single default discount parameter $D=0.75$, as explained in Section 6.3. We set $N=10$ as an order of the model to ensure that most grammatical relations among the tokens of a sentence are captured, even in cases of long sentences. However, the order of the model $N$ could be considered as a hyperparameter of the algorithm (lines 10 and 13 in Algorithm 1).
@@
Assuming the Principle of Linguistic Individuality that at any time $\tau$, for any language $L_{\tau}$, there do not exist two individuals, $\mathcal{A}$ and $\mathcal{B}$, for whom their grammar is identical [97], we expect that in a large number of cases the two Grammar Models $G_{\mathcal{A}}$ and $G_{\mathcal{B}}$ would assign two different probabilities to the same token in context. $P(t\mbox{}_{k}|t\mbox{}_{<k};G_{\mathcal{A}})$ can be interpreted as the extent to which token $t\mbox{}_{k}$ is ’grammatical’ or, according to Cognitive Linguistic theories, entrenched for author $\mathcal{A}$ when in context $t\mbox{}_{<k}$. For this reason, we would expect that, if the sequence $\sigma=(t\mbox{}_{1},t\mbox{}_{2},t\mbox{}_{3})$ was produced by $G_{A}$, then $P(\sigma;G_{\mathcal{A}})>P(\sigma;G_{\mathcal{B}})$ or, more generally according to the Principle of Linguistic Individuality, for any $\mathcal{A}^{\prime}\neq\mathcal{A}$, $P(\sigma;G_{\mathcal{A}})>P(\sigma;G_{\mathcal{A}^{\prime}})$. The model $G_{\mathcal{A}}$ is estimated from $\mathbb{S}_{\mathcal{A}}\mbox{}$ (line 10 in Algorithm 1). To approximate a comparison population of $G_{\mathcal{A}^{\prime}}$, we estimate a set of reference Grammar Models $\mathbb{S}_{\textsf{ref}}\mbox{}=\{G_{1},G_{2},\dots,G_{r}\}$ where each $G_{r}$ is generated by uniformly sampling randomly from $\mathbb{S}_{\textsf{ref}}\mbox{}$ a set of sentences $\mathbb{S}_{r}$ of the same size as the set of sentences in the candidate author’s corpus, $\mathbb{S}_{\mathcal{A}}\mbox{}$, so that $|\mathbb{S}_{\text{ref}}|=|\mathbb{S}_{\mathcal{A}}\mbox{}|$ (line 12 in Algorithm 1). We set $r=100$ but this value can also be considered a hyperparameter of the algorithm. A comparison of hyperparameter settings is reported in Section 4. Taking an approach consistent with the likelihood ratio framework adopted in forensic science, similarly to Ishihara [61], it is therefore possible to calculate (lines 17–19 in Algorithm 1) a mean ratio of the likelihoods assigned to each token in context by $G_{\mathcal{A}}$ and by each $G_{r}$, which we call $\lambda_{G}$:
$\lambda_{G}(t\mbox{}_{k}|t\mbox{}_{<k})=\frac{1}{r}\sum_{j=1}^{r}{\log\frac{P(t\mbox{}_{k}|t\mbox{}_{<k};G_{\mathcal{A}})}{P(t\mbox{}_{k}|t\mbox{}_{<k};G_{j})}}$ (2)
@@
In Eq. 2 the numerator can be seen as a measure of similarity, how much the token is grammatical or entrenched for $G_{\mathcal{A}}$ given its context, while the denominator is a measure of typicality, how atypical this token is in this context in the reference population. A value of $\lambda_{G}$ for each $S\mbox{}_{i}\in\mathbb{S}_{\mathcal{U}}\mbox{}=\{S\mbox{}_{1},S\mbox{}_{2},\dots,S\mbox{}_{m}\}$, can then be calculated (line 15–19 in Algorithm 1) as:
$\lambda_{G}(S\mbox{}_{i})=\sum_{k=1}^{z}\lambda_{G}(t\mbox{}_{k}|t\mbox{}_{<k})$ (3)
@@
Since we model each $S\mbox{}_{i}\in\mathbb{S}_{\mathcal{U}}\mbox{}$ as independent from each other, we can then calculate (line 14–20 in Algorithm 1):
$\begin{split}\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})=\sum_{i=1}^{m}\lambda_{G}(S\mbox{}_{i})=\sum\limits_{i=1}^{m}{\sum\limits_{k=1}^{z}\lambda_{G}(t\mbox{}_{k}|t\mbox{}_{<k})}\end{split}$ (4)
@@
An advantage of this formula is that the final $\lambda_{G}$ can be decomposed into $\lambda_{G}$ values for each single sentence and each single token in each single sentence. This means that the analyst can reconstruct step by step what features contribute to the final score, for example, by ranking each sentence for importance and, for each sentence, color-code or highlight the most important tokens, as demonstrated in Section 7.